13. How many litres of a 12% sulphuric acid solution should be mixed with 5L of a 2% sulphuric acid solution to obtain a 8% solution of sulphuric acid?
1
Expert's answer
2020-04-23T13:10:42-0400

Let "x" litres of a 12% sulphuric acid solution should be mixed with 5L of a 2% sulphuric acid solution to obtain a 8% solution of sulphuric acid.

Concentration of mixture "=8=(12*x+2*5)\/(x+5)"

"\implies 8x+40=12x+10 \implies 3x=30 \implies x=10L"

Thus, 10 litres of a 12% sulphuric acid solution should be mixed with 5L of a 2% sulphuric acid solution to obtain a 8% solution of sulphuric acid.